40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 14:42:34 2011, steve, Update, SUS, oplev gains today 
    Reply  Wed Jul 20 19:24:12 2011, rana, Update, SUS, oplev gains today Untitled.pnga.png
       Reply  Wed Jul 20 19:48:03 2011, Jamie, Update, SUS, Re: oplev gains today 
       Reply  Thu Jul 21 09:04:59 2011, valera, steve, Update, SUS, oplev gains were not optimized 
       Reply  Mon Jul 25 00:35:44 2011, rana, Update, SUS, something wrong with ETMY LR sensor a.pngbad.png
          Reply  Tue Jul 26 11:38:45 2011, steve, Update, SUS, ETMY bias checked 
    Reply  Thu Jul 21 16:06:35 2011, steve, Update, SUS, oplev gains today checked by conlog 
Message ID: 5010     Entry time: Thu Jul 21 09:04:59 2011     In reply to: 5004
Author: valera, steve 
Type: Update 
Category: SUS 
Subject: oplev gains were not optimized 

 

Hi Steve,

 
I did change the ETMY optical lever configuration: http://131.215.115.52:8080/40m/4795
And I left it in that state per Jamie's request.
 
I was going to work on the servo tuning but found that the whitening was not working at that time.
What I was going to do is to measure the open loop gain to make sure the servo is stable, then 
measure the noise and minimize the rms motion by tuning the gain and the filter transfer function.
 
I plan to come to the 40m lab on August 22 for two weeks.
 
Valera.
ELOG V3.1.3-